OpenStudy (anonymous):

the speed of train A is 14 mph slower than the speed of train B. Train A travels 210 miles in the time it takes train B to travel 280 miles. Find the speed of each train.

OpenStudy (bahrom7893):

Va = Vb - 14 Va = 210/t Vb = 280/t

OpenStudy (bahrom7893):

Vb - 14 = 210/t Vb = 280/t

OpenStudy (bahrom7893):

280/t - 14 = 210/t

OpenStudy (bahrom7893):

multiply everything by t: 280 - 14t = 210 -14t = 210 - 280 -14t = -70 t = -70 / -14 = 5

OpenStudy (bahrom7893):

Velocity, or speed of Train A: Va = 210 / 5 = 42mph Train B: Vb = 280 / 5 = 56mph

OpenStudy (anonymous):

thanks for the help.

OpenStudy (anonymous):

A Mathematic solution with comments is attached.